Prev Next

SW&IT Meeting Minutes 2019-11-11

Attendees: Andrew, Carl, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20191104
LR20191104 - Bec, Matthieu 40.0
LR20191104 - Holmberg, Carl 0.0
LR20191104 - Colson, Andrew 0.0
LR20191104 - Edmeades, David 40.0
LR20191104 - Sola, Igor 0.0
LR20191104 - Kubanek, Petr 50.0
LR20191104 - Hooper, Stephen 80.0
LR20191104 - Zhang, Yang 38.0
LR20191104 - Zhang, Xianyu 0.0

News from management

  • Veterans day, no meeting

Previous Action Items

Open topics

Team reports

DE

  • Reset colo project
  • Research possible network workarounds to appease NetOps
  • PEPSI final hardware configs
  • Mountain Synology data move
    • Prep for proposed space upgrade of volume2
    • Order test disks?

MB

  • OC, WPC, Management
  • SOUL, LBTI, Archive, meetings
  • DMS clean up
  • follow-up mt-archive SSD discussion
next
  • Restart checkout list
  • PEPSI wrap up
  • CentOS stream

PK

  • Eigen mean and sum problem
  • newGCS: AZCam.cpp progress, azcamclient, more imagetest, FITSpp.hpp documents
  • IT7640 fix - PR 115
  • TCS build 2019B installed and running
  • support during engineering
  • patches to 2019B - IT7640, IT5140 fixes (elevation re-setting, unwrap calls being trigerred by instrument authorization)
  • TCS README.md

YZ

  • Worked with Olga to test the MODS GUI problem (IT#7855). MODS GUI has problem with only one menu updating in the offset GUI. To be investigated further.
  • Patched the TCS IIF instrument C interface (pull request merged to TCS master branch).
  • Continue to study and work on the fix for the LBC non-sidereal guiding issue (IT#7847).
  • Discussed with Al (together with Xianyu) on several AO-Cravat issues (Strehl and FWHM calculation, OSCO metadata etc). Al promised to organize a meeting to discuss plans on these issues.
  • Redesigning and rebuilding a new database for AO-Cravat OSCO data, as well as changing the web interface to match the new design.

AC

  • new_GCS: copied PK's rebase and added changes for OAC telemetry
  • argos-dx navigation menu organization
  • PMS is on and collecting data but not saving to telemetry - perhaps related to moving to new host?? Doubtful but will investigate

CH

Last week:
  • ALH:
    • Cleaned up TMS status check script.
    • Coded prototype a mountain UPS and chiller status check script, querying the local Zabbix instance.
    • Provided LUCI data dictionary list to PEPSI team, as an example of what we want from the system.
    • Reviewed Statserv/Indi link methodology to provide background for possible LBTI alert fix.
    • Began reviewing Google Mail API for method to retrieve mountain PLC alert emails. Made initial attempt to create system to system authentication cert.
  • SupportScripts: Pushed code changes to github.

This week:
  • ALH: Prototype a Google Mail API-based script to retrieve mountain PLC alert emails.

XZ

Last week:
  • SOUL commission night observation (Wednesday-Friday)
  • Improve the jumping actuator analyze routine
  • Transfering files for jumping actuator analyze
  • Write a rountine push jumping actuator analyze data structure to MongoDB

This week:
  • SPIE paper abstracts
  • TNO ASM meeting
  • conitune AO-data document
  • jumping actuator analyze demonstration
  • Setup the mountain anallyze VM

IS

Last week:
  • miniQ
    • Kept working on miniQ version. Was use successfully on Monday but still needs some work.
  • PIT
    • Discussed with Michelle ways to improve and automatize proposals after period closes.
    • Discussed/plan to add a resubmit option to the PIT so there are no or to limit duplicates
This week
  • miniQ
    • Still have to work on the report part, and routing and history navigation.
  • PIT
    • work on ways to improve and automatize proposals after period closes.
    • work on implementing a resubmit option to the PIT so there are no or to limit duplicates

SH

  • mountain Tuesday
    • add second interface to VM1
    • cleanup ethernet in CRA
    • new ethernet connector in control room
    • 4TB drive installation in old VM1
  • forensics on archive/TCS issues
  • ao-analysis completed
  • rework mountain cr-ups to be in zabbix
  • moved some VMs off of Synology
Topic revision: r12 - 18 Nov 2019, StephenHooper
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback